OpenStudy (anonymous):

which way does this parabola open: x^2 = -8(y-3)

OpenStudy (anonymous):

downward.

OpenStudy (anonymous):

Down, because the coefficient of \(-x^2\) ends up being negative.

OpenStudy (anonymous):

thank you! but can someone show me how to figure that out? I am completely lost :(

OpenStudy (anonymous):

You just get it in the form \(ax^2 + bx + c = 0\) (They don't have to be +'s, any of those can be negative)

OpenStudy (anonymous):

Then if a is positive, it opens up, and if a is negative, it opens down.

OpenStudy (anonymous):

In this case, you had to divide by -8 which caused a negative to show up. So it opens down.

OpenStudy (anonymous):

how to you plot x^2=y? since x is the independent variable, for different values of x gives out different values of y. so, if we take all the ordered pairs (x,y) and try plotting theses points on the y-vs-x graph, you would get a concave upward graph, with it's lowest point (0,0). This curve is called a parabola. and since it opens upward, it is called upward parabola. If we now try to graph y=-x^2, we get a downward parabola.

OpenStudy (anonymous):

so I am trying to get all #'s to the left of the equal sign correct? That is started by dividing both sides by -8? Or do I complete the : -8(y-3) first? I have this formula: x=(y-k)^2 + h but I don't know how to drop in the values with information I have been given. thank you !!!!

OpenStudy (anonymous):

always try to know how we do rather than what we do.

OpenStudy (anonymous):

Wait, are you trying to complete the square? Because you don't need to know that just to know which way the parabola opens. All you have to do to know that is solve for y.

OpenStudy (anonymous):

no I am not trying to complete the square.. I just don't know how to find the (h,k), focus, radius, stuff like that

OpenStudy (anonymous):

Oh! I get it now. Sorry.

OpenStudy (anonymous):

Do you know \(4p(y-k) = (x-h)^2\) That's basically the form the equation given was in.

OpenStudy (anonymous):

i did not know that ! is that the best formula to attack these questions with?

OpenStudy (anonymous):

Well especially when they give you the question in that form. And it gives you h and k right there. To put it in that form when not given like that, you'll usually have to complete the square.

OpenStudy (anonymous):

thank you so much! i guess i need to brush up on the completing the square :)

OpenStudy (anonymous):

I can help you with that, too. Do you know the quadratic formula?

OpenStudy (anonymous):

just about the only formula I do know :) so yes :)

OpenStudy (anonymous):

You're like me, then. The good news is you can use the quadratic formula to complete the square.

OpenStudy (anonymous):

Start with the classic \[x = \frac{-b \pm \sqrt{(b^2 - 4ac)}}{2a}\]

OpenStudy (anonymous):

\[x + \frac{b}{2a} = \frac{\sqrt{(b^2 - 4ac)}}{2a}\] \[(x + \frac{b}{2a})^2 = \frac{(b^2 - 4ac)}{4a^2}\]

OpenStudy (anonymous):

Look familiar? :D

OpenStudy (anonymous):

the top one did :)

OpenStudy (anonymous):

The quadratic formula is derived by completing the square, I just reversed it a little to show how it can help

OpenStudy (anonymous):

i see.. but not really :D math is my nemesis!! lol..

OpenStudy (anonymous):

Ok, well, you might have to brush up on it then. :P

OpenStudy (anonymous):

I can't really teach HOW to do it... just how to do it faster. :D

OpenStudy (anonymous):

hahah.. i appreciate all your help and i will see if i can figure this out :)

OpenStudy (anonymous):

lol